Grade 9General Physics

two racing cars of mass m1 and m2 are moving in cicles of radii r1 and r2 respectively.their speeds are such that each makes a complete circle in the same lengthof time t .the ratio of the angular speed of the first car to that of the second car is ?

garima singh
10 Years agoGrade 9

1 Answer

Subhiksha
10 Years ago
1:1
Because angular speed = linear velocity/ radius
Here, v =
 w = v/r= 2(pi)r/tr
w= 2(pi)/t
 
Here, since t is equal in both cases, w1:w2 = 1;1